OpenStudy (anonymous):

please help! write an equation of the line that has a slope of -3 and contains the point (1,1) in point-slope form

OpenStudy (anonymous):

i dont know how to write the equation in slope -intercept form

OpenStudy (ahsome):

Point slope form is written in this form\[y=mx+b\]Where: \(y\) and \(x\) is the y and x value \(m\) Is the slope \(b\) is just a number, and is equal to the y-intercept. In your case, we are given 3 things, the y and x value (1,1), and the slope, -3. Now, we need to find \(b\). What we can do is simply put all the numbers we know into the equation, and find \(b\)

OpenStudy (ahsome):

\[y=mx+b\]\[1=6\times 1+b\]\[1=6+b\]\[1-6=b\]\[-5=b\]\[b=-5\]

OpenStudy (ahsome):

Now, pretty much put all the numbers into the equation \(y=mx+b\) \(y=6x-5\) Do NOT put the x and y values you have (1,1), since they can change anytime when given a different point. All that stays the same is the slope, m, and the y-intercept, b.

OpenStudy (ahsome):

That should help, @priea

OpenStudy (anonymous):

so what is the equation i should put down because it said to just write the equation @Ahsome

OpenStudy (ahsome):

\(y=6x−5\)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@Ahsome

OpenStudy (ahsome):

Yeah?

OpenStudy (anonymous):

the answer is y=6x-5?

OpenStudy (ahsome):

Yah. Unless I did my calculations wrong. Will check...

OpenStudy (ahsome):

Whoops, I put 6 for m, not -3.

OpenStudy (ahsome):

\(y=mx+b\) \(1=-3×1+b\) \(1=-3+b\) \(4=b\) \(b=4\) So the answer is \(y=-3x+4\)

OpenStudy (anonymous):

okay thanks ill put that in and see. but ill probably need more help. LOL

OpenStudy (ahsome):

Sure :)

OpenStudy (anonymous):

write an equation of the line that has a sloe of 3 and contains the point (2,5) in slope-intercept form

OpenStudy (anonymous):

*slope

OpenStudy (ahsome):

Do the same thing I did before, but just change the slope and point to 3 and (2,5)

OpenStudy (anonymous):

but i dont want to solve it i just need an equation